OpenStudy (insert_username):

6(2x - 11) + 15 = 21 Write the steps you will use to solve the equation and explain each step. I'm having some trouble with writing it out step by step. I know that x= 6, but I don't know how to write and explain the steps since I did it in my head.

OpenStudy (anonymous):

\[6(2x-11)+15=21\] Subtract the 15 over\[6(2x-11)=6\] Distribute the 6\[12x-66=6\]Add the 66 over\[12x=72\] Divide the 12 over to get the x alone\[x=6\] This is how I would write it out step by step. You don't really need to include the explanations, your teacher would understand if you just put down the steps.

OpenStudy (insert_username):

Thank you so much for helping. Usually I just do the problems in my head so I'm not that good at writing down the steps, so thanks.

OpenStudy (anonymous):

Yeah, I understand, I used to do all of these kinds of problems in your head. Once you get into the higher levels of math, you have to write stuff down though, since it's near impossible to do in your head :)

OpenStudy (insert_username):

Thanks! I'm gonna fan you